# Constructing Variables Using Classifiers as an Aid to Regression: An Empirical Assessment

### 11 Mar 2024 v1

@@ 1 Introduction
In the case of regression, this involves predicting the value of a numerical variable (noted $y$), for example the amount of an invoice, using a set of explanatory variables (a vector noted $X$). In the case of machine learning, we seek to learn a function $f$ such that $y=f(X)$ using a machine learning algorithm and a learning set, a set of $N$ input-output pairs $(X_{i},y_{i}),i={1,...,N}$. During this modeling stage, there is often a need to create new variables that better describe the problem and enable the model to achieve better performance. This is what we call the "engineering process of creating new explanatory variables" [19]. In this case, the idea is that the new variables (a vector, in this case $X^{\prime}$) will provide additional information. By automating the generation of these "new variables", more useful and meaningful information can be extracted from the data, in a framework that can be applied to any problem. This allows the machine learning engineer to devote more time to more useful tasks.
The aim of the article presented here is to propose a method for automatically creating variables (in the case of regression) that complement the information contained in the vector $X$ to predict the values of the dependent variable $y$. The proposed method first transforms the regression problem into several classification sub-problems, then integrates the results in the form of additive variables ($X^{\prime}$). The ’augmented’ vector, $X^{\prime\prime}=X\cup X^{\prime}$, is then fed into standard regressors to measure the contribution of the variables created. The benefits of this approach are presented in the form of a detailed experimental study.
@@ 2 Proposal
The proposed method relies on transforming the regression problem into a classification problem to construct the additional variables. This idea has already been suggested, but our approach differs from the state of the art, as explained below in section 2.1. This difference having been presented, section 2.2 describes in detail the proposed method and section 2.3 the choices that have been made for its implementation.
@@ 2.1 Related work
Solving a regression problem using classification models is an approach that has already been explored. This process has been described in numerous papers [2, 3, 14, 15, 17, 20] and generally consists of two main steps: (i) discretization (also called partitioning or binning) of the target variable to enable the use of classifiers on the dataset; (ii) prediction of the regression is then usually performed by calculating the mean ([2, 3, 14, 17]) or median ([3, 15, 20]) of the instances within the fragment of discretized output that the classifier has predicted.
The method we propose and present below differs from these works in that the classifiers used by the method are solely aimed at adding complementary explanatory variables to the initial explanatory variables (native variables). The augmented vector is then positioned at the input of a standard regressor. This regressor directly predicts the target variable without any further transformation or estimation operations. As the next section will show, the proposed method is linked to a conditional estimate of the density function of $y$. It would then be interesting to test other, less computationally-intensive methods of variable creation in this same framework [18, 13, 21, 16] in future work. But the general principle remains the same.
@@ 2.2 General principle
Once the classes have been encoded and one (or more) classifiers trained (using the training set), it is then possible to predict the membership of individuals in the previously defined classes. The predictions of the classifier(s) on each individual will then be used to create a new "extended" data set, whether for the training set or the test set.
Figure 2 illustrates the extension from vector $X$ (with $d$ components) to vector $X^{\prime\prime}$ (with $d+S$ components). Section 2.3 will describe in more detail what the vector $X^{\prime}=\{X^{\prime}_{1},...,X^{\prime}_{i},...,X^{\prime}_{S}\}$ from the prediction of the classifier(s) is composed of.
@@
Once the $X$ vector has been "extended", the regression model can be trained and predicted on the new data. We can then compare the regressor’s performance when trained on the original data set (with only $X$) with its performance when using the extended set ($X^{\prime\prime}=X\cup X^{\prime}$). The assumption being that the regression model will have a better intuition of the general position of individuals in the space of the variable to be regressed and will present better results.
Figure 3 summarizes the process followed by the proposed method: (i) the first step is to transform the regression problem into a classification problem. To do this, the target variable $y$ is first discretized (see section 2.3), then classes are defined using the thresholds thus defined. (ii) In a second step, the classifiers are trained using the original $X$ descriptive variables and the new classes derived from $y$. Then the prediction, of the classifiers using the initial $X$ vector, is used to extract new features, i.e. $X^{\prime}$. (iii) Finally, the regression model can be trained using the vector $X^{\prime\prime}=X\cup X^{\prime}$. The proposed method is therefore based on a discretization and class encoding mechanism. These two processes can be implemented in different ways. An example of implementation is given below in section 2.3.
@@ 2.3 Implementation
One of the most important aspects of the proposed method is the association of classes with instances. These classes will be used to train the classifiers. This process is based on two steps: the definition of thresholds and the encoding of classes. For the definition of threshold placement, which amounts to an unsupervised discretization of the variable to be regressed, there are numerous possibilities in the literature, such as “EqualWidth”, “EqualFreq”, etc. For class encoding, it is possible to pose the problem either as an $S$ class classification problem (where $S$ denotes the number of thresholds), or as $S$ binary classification problems, or as a multi-label classification problem.
The choice of discretization method and the number of thresholds are linked, and this choice represents a compromise between (i) classification efficiency and (ii) information input for the initial regression problem. In the case of a single classifier that predicts the membership interval of individuals, it seems obvious that the larger S is, the more precise the information provided to the regressor will be, but the more difficult the problem will become to learn. On the other hand, if we pose the problem as $S$ binary classification problems, each classifier will have a problem of the same difficulty to solve, regardless of the value of $S$. Preliminary tests carried out on the set of datasets described in Section 3.1 have verified this behavior. The final choices for the various elements associated with the method proposed in this article are described below.
For the discretization method - The “Equal Frequency" method was chosen because, unlike "Equal Width", it does not risk creating intervals that contain no individuals. It also ensures that there are no classification problems where the minority class (extreme threshold on the left, see Figure 1) represents less than $\frac{1}{S+1}$ percent of individuals.
@@
For the number of thresholds - The first intuition (which was confirmed during experimentation) is that the greater the number of thresholds defined, the greater the potential performance gain. Indeed, the more thresholds defined, the finer the discretization and the closer the classifiers’ prediction will be to the true regression value. In the experimental phase of this article, we will illustrate the performance of our method as a function of the number of thresholds defined. The number of thresholds defined on the space of the variable to be regressed is therefore a hyper-parameter of our method.
For the definition of classes associated with thresholds - We chose classes linked to value inferiority thresholds ($C_{i}:=\mathbb{1}_{y\leq y_{i}}$) (see Section 2.2 and Figure 1).
For the variables extracted from the classifiers - Each classifier will be trained to predict whether the data provided is below or above the threshold at which it is assigned a class. As the method is intended to be generic, it was decided to extract the conditional probabilities predicted by each classifier. Indeed, this information is available for the vast majority of classifiers in the literature. Conditional probabilities [^1] Note: as the classifiers selected are random forests, the conditional probability corresponds to the proportion of trees that voted for the class in question. of class 1 (i.e. if $y\leq y_{i}$) predicted by each classifier will therefore make up the vector $X^{\prime}=\{X^{\prime}_{1},....,X^{\prime}_{i},...,X^{\prime}_{S}\}=\{P(C_{1}=1|X),...,P(C_{i}=1|X),...,P(C_{S}=1|X)\}$; $X^{\prime}$ will then be added to the initial data vector ($X$).
@@ 3.1 Data sets and pre-processing
To carry out the analysis of the proposed method, we selected a large collection of regression datasets from the UCI Repository [10] or from Kaggle [1]. A total of 33 datasets were used, of which 23 consisted of more than 10,000 individuals. The remaining 10 datasets range from 1,030 to 9,568 individuals (see Table 1). This selection was largely influenced by the article by M. Fernandez-Delgado et al. [11]. The datasets chosen were, according to the categorization of Fernandez-Delgado et al. "large and difficult". Some of these datasets included several subsets of data with different target variables to regress, and therefore different regression problems. It was chosen not to include several regression problems from the same datasets, so as not to give more weight to certain datasets and create a bias in the comparison of methods that might favor one method over another. The reader can refer to the "Dataset name" column of the Table 1, which designates the regression problem used for each dataset.
@@
Train-Test split and model optimization: Each dataset has been split into a "10-fold cross validation", giving 10 training and test sets. For the three models requiring parameter optimization (AR, FA and XGB see Section 3.3): 30% of the training set was reserved for optimizing model parameters in a "grid-search" process, often avoiding over-training. Then, provided with the "right" training parameters, the model was trained using the remaining 70% of the training set. Finally, once the model had been learned for this fold, its performance was measured. This rigorous process was carried out for all datasets, all folds and all $S$ values, resulting in the training of thousands of models, but enabling a rigorous test of the proposed method.
@@ 3.2 Metric used for results
When comparing regression models, the aim is to find a model without over- or under-fitting that achieves a low generalization error, which characterizes its predictive performance. Various metrics are proposed in the literature, such as the root mean square error (RMSE), the mean absolute error (MAE), the R-Square, etc… There does not seem to be a consensus on the "best" metric to use, although some papers do offer comparisons [5]. In the rest of the results, RMSE has been chosen and is defined as: $RMSE=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}}$, where $n$ is the number of individuals, $y_{i}$ the desired value and $\hat{y}_{i}$ the output of a regressor.
@@ 3.3 Tested Regressors
During the experiments, five regressors using different frameworks were used. They are described briefly below, although they are widely known in the machine learning community.
@@
$\bullet$ Linear Regression (LR): Regression refers to the process of estimating a continuous numerical variable using other variables correlated with it. This means that regression models take the form of $y=f(X)$ where $y$ can take on an infinite set of values. In this article, we have chosen a multivariate linear regression whose model, $f$, is determined by the method of least squares. We used Scikit-Learn version 0.24.2. This model does not require parameter optimization.
@@
$\bullet$ Regression Tree (DT): Decision trees are used to predict a real quantity which, in the case of regression, is a numerical value. Algorithms for building decision trees are usually constructed by dividing the tree from the root to the leaves, choosing at each step an input variable that achieves the best sharing of the set of individuals. In the case of regression trees, the aim is to maximize inter-class variance (i.e. to have subsets whose values of the target variable are as widely dispersed as possible). In this article, we used Scikit-Learn version 0.24.2. This model requires parameter optimization.
@@
$\bullet$ Random Forest (RF): Decision-tree forests (or Random Forest (RF) regressors) were first proposed by Ho in 1995 [12] and then extended by Leo Breiman and Adele Cutler [8]. This algorithm combines the concepts of random subspaces and bagging. The decision tree forest algorithm performs training on multiple decision trees trained on slightly different data subsets. In this article, we used Scikit-Learn version 0.24.2. This model requires parameter optimization.
@@
$\bullet$ XGBoost (XGB): XGBoost [9] is a boosting method. It sequentially combines weak learners that would individually perform poorly to improve the prediction of the full algorithm. A weak learner is a regressor with poor regression performance. In this boosting algorithm, high weights will be associated with weak learners having good accuracy, and conversely, lower weights with weak learners having poor accuracy. In the training phase, high weights are associated with data that has been poorly "learned", so that the next weak learner in the sequence will focus more on this data. In this article, we used XGBoost version 1.4.1. This model requires parameter optimization.
@@
$\bullet$ Selective Naïve Bayes (SNB): In the context of regression, a naive Bayes regressor (NB) whose variables are weighted by weights (NBP) can be obtained in two steps. First, for each explanatory variable, a 2D grid is created to estimate $P(X,y)$ (see for example [6]). Then, in a second step, all variables are grouped together in a Forward Backward algorithm [7] to estimate their informativeness in the context of a Naive Bayes regressor. At the end of the second step, the final model (to be deployed) is a Naïve Bayes (which uses the 2D discretization found in the first step) where the variables are weighted (the weights are found in the second step). In this article we used the Weighted Naïve Bayes produced by Khiops library (www.khiops.org). This model does not require parameter optimization.
@@ 4 Results
$\bullet$ Illustrative results - Figure 4 illustrates the behavior of the proposed method: on the left for linear regression and the "KEGG Metabolic Reaction" dataset, in the middle and on the right for the "SML 2010" dataset, respectively for the Selective Naïve Bayes regressor and the Random Forest. These 3 figures are fairly representative of the results obtained, with a more or less marked decrease in RMSE versus the value of $S$ and the regressor considered. This reduction is most marked at the start of the curve and then levels off for a less marked gain beyond $S=16$. The following section presents the results obtained for each dataset.
@@
$\bullet$ Results tables on all datasets - The results obtained are presented in detail in Table 2 for a number of thresholds equal to 32 ($S=32$). For each dataset, this table gives the test results (average results over the 10 test folds (see Section 3.1)) for each of the five regressors for which the addition of the $X^{\prime}$ vector was tested. In this table, a value in bold indicates a significant difference between results with the $X$ vector "Native" and the $X^{\prime}$ vector "Aug" (Increased) according to paired Student’s t test (p-value at 5%). The last line of the table shows the number of losses, ties and wins for "Aug" versus "Native". We can see that the addition of the $X^{\prime}$ vector essentially benefits 3 of the regressors: the linear regression, the regression tree and the naive Bayes regressor. For the random forest and XGboost, the gains and/or losses are fairly limited.
@@
The penultimate row of the table shows the average RMSE obtained on all data sets (purely indicative), confirming the interest of the proposed method for three of the five regressors. This average should be treated with caution. As each dataset is a problem of unique difficulty, the error scale differs between each dataset. So, if the new RMSE mean in test is higher, this does not necessarily mean that the regressors perform worse on average. For this reason, in the next section we present another view of the results for further analysis.
$\bullet$ Critical diagram - We present in Figure 5 critical diagrams comparing the results obtained by each regressor through the Nemenyi post-hoc test, performed after a Friedman signed rank test of RMSE values, considering all data sets.
@@
It can be seen from the diagram on the left of figure 5 that the XGBoost regressor shows the best performance, obtaining the highest mean RMSE rank, closely followed by random forest (RF). These two regressors outperform the other three regressors by a wide margin, with linear regression being the worst-performing model. It is closely followed by the naive Bayes regressor and the regression tree. The figure 5 in the center compares the performance of all regressors, with and without the proposed method: the first point that we can observe is that, statistically, all regressors with $X^{\prime}$ have a better average rank than their alter-ego with only $X$. This is an encouraging result: no regressor has been negatively affected by the proposed method, since all regressors who have used this method are better ranked than their basic versions. Different groups are again present, and the reader can compare the change in ranking of different regressors by comparing the figure on the left with the one in the middle. Finally, the diagram on the right of figure 5 shows that: when the proposed method is used, the regressors are no longer differentiable (only one group compared to the left figure) according to the Nemenyi test, even though they have different mean ranks.

### 13 Mar 2024 v2

@@ 1 Introduction
In the case of regression, this involves predicting the value of a numerical variable (noted $y$), for example the amount of an invoice, using a set of explanatory variables (a vector noted $X$). In the case of machine learning, we seek to learn a function $f$ such that $y=f(X)$ using a machine learning algorithm and a training set, a set of $N$ input-output pairs $(X_{i},y_{i}),i={1,...,N}$. During this modeling phase, it is often necessary to create new variables that better describe the problem and enable the model to achieve better performance. This is what we call the "engineering process of creating new explanatory variables" [19]. The idea here is here that the new variables (a vector, in this case $X^{\prime}$) will provide additional information. By automating the generation of these "new variables", more useful and meaningful information can be extracted from the data, in a framework that can be applied to any problem. This allows the machine learning engineer to devote more time on more useful tasks.
The aim of the article presented here is to propose a method for automatically creating variables (in the case of regression) that complement the information contained in the vector $X$ in order to predict the values of the dependent variable $y$. The proposed method first transforms the regression problem into several classification sub-problems, then integrates the results in the form of additive variables ($X^{\prime}$). The ’augmented’ vector, $X^{\prime\prime}=X\cup X^{\prime}$, is then fed into standard regressors to measure the contribution of the created variables. The advantages of this approach are presented in the form of a detailed experimental study.
@@ 2 Proposal
The proposed method relies on transforming the regression problem into a classification problem to construct the additional variables. This idea has been proposed before, but our approach differs from the state of the art, as explained below in section 2.1. Having presented this difference, section 2.2 describes the proposed method in detail and section 2.3 describes the choices made for its implementation.
@@ 2.1 Related work
Solving a regression problem using classification models is an approach that has already been explored. This process has been described in numerous papers [2, 3, 14, 15, 17, 20] and generally consists of two main steps: (i) discretization (also called partitioning or binning) of the target variable to enable the use of classifiers on the dataset; (ii) prediction of the regression is then usually performed by calculating the mean ([2, 3, 14, 17]) or median ([3, 15, 20]) of the instances within the fragment of discretized output predicted by the classifier.
The method we propose and present below differs from these works in that the classifiers used by the method are solely aimed at adding complementary explanatory variables to the initial explanatory variables (native variables). The augmented vector is then positioned at the input of a standard regressor. This regressor directly predicts the target variable without any further transformation or estimation operations. As the next section will show, the proposed method is linked to a conditional estimate of the density function of $y$. It would then be interesting to test other, less computationally-intensive methods of variable creation in the same framework [18, 13, 21, 16] in future work. But the general principle remains the same.
@@ 2.2 General principle
Once the classes have been encoded and one (or more) classifiers have been trained (using the training set), it is then possible to predict the membership of individuals in the previously defined classes. The predictions of the classifier(s) on each individual will then be used to create a new "extended" data set, either for the training set or the test set.
The figure 2 illustrates the extension of the vector $X$ (with $d$ components) to the vector $X^{\prime\prime}$ (with $d+S$ components). The section 2.3 will describe in more detail what the vector $X^{\prime}=\{X^{\prime}_{1},...,X^{\prime}_{i},...,X^{\prime}_{S}\}$ from the prediction of the classifier(s) is composed of.
@@
Once the $X$ vector has been "extended", the regression model can be trained and predicted on the new data. We can then compare the regressor’s performance when trained on the original data set (using only $X$) with its performance when using the extended set ($X^{\prime\prime}=X\cup X^{\prime}$). The assumption being that the regression model will have a better intuition of the general position of individuals in the space of the variable to be regressed and will produce better results.
Figure 3 summarizes the process followed by the proposed method: (i) the first step is to transform the regression problem into a classification problem. To do this, the target variable $y$ is first discretized (see section 2.3), then classes are defined using the thresholds thus defined. (ii) In a second step, the classifiers are trained using the original descriptive variables $X$ and the new classes derived from $y$. Then the prediction, of the classifiers using the initial vector $X$, is used to extract new features, i.e. $X^{\prime}$. (iii) Finally, the regression model can be trained using the vector $X^{\prime\prime}=X\cup X^{\prime}$. The proposed method is thus based on a discretization and a class encoding mechanism. These two processes can be implemented in different ways. An example of implementation is given below in the section 2.3.
@@ 2.3 Implementation
One of the most important aspects of the proposed method is the association of classes with instances. These classes will be used to train the classifiers. This process is based on two steps: the definition of thresholds and the encoding of classes. For the definition of threshold placement, which amounts to an unsupervised discretization of the variable to be regressed, there are numerous possibilities in the literature, such as “EqualWidth”, “EqualFreq”, etc. For class encoding, it is possible to formulate the problem either as an $S$ class classification problem (where $S$ denotes the number of thresholds), or as $S$ binary classification problems, or as a multi-label classification problem.
The choice of discretization method and the number of thresholds are linked, and this choice represents a trade-off between (i) classification efficiency and (ii) information input for the initial regression problem. In the case of a single classifier predicting the membership interval of individuals, it seems obvious that the larger S is, the more precise the information provided to the regressor will be, but the more difficult the problem will become to learn. On the other hand, if we formulate the problem as $S$ binary classification problems, each classifier will have a problem of the same difficulty to solve, regardless of the value of $S$. Preliminary tests carried out on the set of datasets described in Section 3.1 have confirmed this behavior. The final choices for the various elements associated with the method proposed in this article are described below.
For the discretization method - The “Equal Frequency" method was chosen because, unlike the “Equal Width” method, it does not run the risk of creating intervals that contain no individuals. It also ensures that there are no classification problems where the minority class (extreme threshold on the left, see Figure 1) represents less than $\frac{1}{S+1}$ percent of the individuals.
@@
For the number of thresholds - The first intuition (confirmed by the experiments) is that the greater the number of thresholds defined, the greater the potential performance gain. Indeed, the more thresholds defined, the finer the discretization and the closer the classifiers’ prediction will be to the true regression value. In the experimental phase of this article, we will illustrate the performance of our method as a function of the number of thresholds defined. The number of thresholds defined on the space of the variable to be regressed is therefore a hyper-parameter of our method.
For the definition of classes associated with thresholds - We chose classes associated with value inferiority thresholds ($C_{i}:=\mathbb{1}_{y\leq y_{i}}$) (see Section 2.2 and Figure 1).
For the variables extracted from the classifiers - Each classifier will be trained to predict whether the data provided is below or above the threshold at which it is assigned a class. As the method is intended to be generic, it was decided to extract the conditional probabilities predicted by each classifier. Indeed, this information is available for the vast majority of classifiers in the literature. Conditional probabilities [^1] Note: as the selected classifiers are random forests, the conditional probability corresponds to the proportion of trees that voted for the class in question. of class 1 (i.e. if $y\leq y_{i}$) predicted by each classifier will therefore make up the vector $X^{\prime}=\{X^{\prime}_{1},....,X^{\prime}_{i},...,X^{\prime}_{S}\}=\{P(C_{1}=1|X),...,P(C_{i}=1|X),...,P(C_{S}=1|X)\}$; $X^{\prime}$ will then be added to the initial data vector ($X$).
@@ 3.1 Data sets and pre-processing
To carry out the analysis of the proposed method, we selected a large collection of regression datasets from the UCI Repository [10] or from Kaggle [1]. A total of 33 datasets were used, of which 23 consisted of more than 10,000 individuals. The remaining 10 datasets range from 1,030 to 9,568 individuals (see Table 1). This selection was largely influenced by the article by M. Fernandez-Delgado et al. [11]. The selected datasets were, according to the categorization of Fernandez-Delgado et al. "large and difficult". Some of these datasets included several subsets of data with different target variables to be regressed, and therefore different regression problems. It was decided not to include multiple regression problems from the same datasets in order to avoid giving more weight to certain datasets and creating a bias in the comparison of methods that might favour one method over another. The reader can refer to the "Dataset name" column of the table 1, which identifies the regression problem used for each dataset.
@@
Train-Test split and model optimization: Each dataset has been split into a "10-fold cross validation", resulting in 10 training and test sets. For the three models requiring parameter optimization (AR, FA and XGB see Section 3.3): 30% of the training set was reserved for optimizing the model parameters in a "grid-search" process, often avoiding over-training. Then, provided with the "right" training parameters, the model was trained using the remaining 70% of the training set. Finally, once the model had been trained for this fold, its performance was measured. This rigorous process was carried out for all datasets, all folds and all $S$ values, resulting in the training of thousands of models, but allowing a rigorous test of the proposed method.
@@ 3.2 Metric used for results
When comparing regression models, the aim is to find a model without over- or under-fitting that achieves a low generalization error, which characterizes its predictive performance. Various metrics are proposed in the literature, such as the root mean square error (RMSE), the mean absolute error (MAE), the R-Square, etc… There seems to be no consensus on the "best" metric to use, although some papers do offer comparisons [5]. In the rest of the results, RMSE has been chosen and is defined as: $RMSE=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}}$, where $n$ is the number of individuals, $y_{i}$ the desired value and $\hat{y}_{i}$ the output of a regressor.
@@ 3.3 Tested Regressors
Five regressors were used in the experiments, using different frameworks. They are briefly described below, although they are well known in the machine learning community.
@@
$\bullet$ Linear Regression (LR): Regression refers to the process of estimating a continuous numerical variable using other variables that are correlated with it. This means that regression models take the form of $y=f(X)$ where $y$ can take on an infinite number of values. In this article, we have chosen a multivariate linear regression whose model, $f$, is determined by the method of least squares. We have used Scikit-Learn version 0.24.2. This model does not require parameter optimization.
@@
$\bullet$ Regression Tree (DT): Decision trees are used to predict a real quantity which, in the case of regression, is a numerical value. Algorithms for building decision trees are usually constructed by dividing the tree from the root to the leaves, choosing at each step an input variable that achieves the best partitioning of the set of individuals. In the case of regression trees, the aim is to maximize the inter-class variance (i.e. to have subsets whose values of the target variable are as widely spread as possible). In this article, we used Scikit-Learn version 0.24.2. This model requires parameter optimization.
@@
$\bullet$ Random Forest (RF): Decision-tree forests (or Random Forest (RF) regressors) were first proposed by Ho in 1995 [12] and then extended by Leo Breiman and Adele Cutler [8]. This algorithm combines the concepts of random subspaces and bagging. The decision tree forest algorithm performs training on multiple decision trees trained on slightly different subsets of data. In this article, we have used Scikit-Learn version 0.24.2. This model requires parameter optimization.
@@
$\bullet$ XGBoost (XGB): XGBoost [9] is a boosting method. It sequentially combines weak learners that would individually perform poorly to improve the prediction of the full algorithm. A weak learner is a regressor with poor regression performance. In this boosting algorithm, high weights will be associated with weak learners having good accuracy, and conversely, lower weights with weak learners having poor accuracy. In the training phase, high weights are associated with data that has been poorly "learned", so that the next weak learner in the sequence will focus more on that data. In this article, we have used XGBoost version 1.4.1. This model requires parameter optimization.
@@
$\bullet$ Selective Naïve Bayes (SNB): In the context of regression, a naive Bayes regressor (NB) whose variables are weighted by weights (NBP) can be obtained in two steps. First, for each explanatory variable, a 2D grid is created to estimate $P(X,y)$ (see for example [6]). Then, in a second step, all variables are grouped together in a Forward Backward algorithm [7] to estimate their informativeness in the context of a Naive Bayes regressor. At the end of the second step, the final model (to be deployed) is a Naïve Bayes (using the 2D discretization found in the first step) where the variables are weighted (the weights are found in the second step). In this article we have used the Weighted Naïve Bayes produced by Khiops library (www.khiops.org). This model does not require parameter optimization.
@@ 4 Results
$\bullet$ Illustrative results - Figure 4 illustrates the behavior of the proposed method: on the left for linear regression and the "KEGG Metabolic Reaction" dataset, in the middle and on the right for the "SML 2010" dataset, respectively for the Selective Naïve Bayes regressor and the Random Forest. These 3 figures are fairly representative of the results obtained, with a more or less marked decrease in the RMSE versus the value of $S$ and the regressor considered. This decrease is most pronounced at the beginning of the curve and then levels off for a less pronounced increase beyond $S=16$. The following section presents the results obtained for each data set.
@@
$\bullet$ Results tables on all datasets - The results obtained are presented in detail in Table 2 for a number of thresholds equal to 32 ($S=32$). For each dataset, this table gives the test results (average results over the 10 test folds (see Section 3.1)) for each of the five regressors for which the addition of the vector $X^{\prime}$ was tested. In this table, a value in bold indicates a significant difference between results with the $X$ vector "Native" and the $X^{\prime}$ vector "Aug" (Increased) according to paired Student’s t test (p-value at 5%). The last row of the table shows the number of losses, ties and wins for "Aug" versus "Native". We can see that the addition of the $X^{\prime}$ vector essentially benefits 3 of the regressors: the linear regression, the regression tree and the naive Bayes regressor. For the random forest and XGboost, the gains and/or losses are fairly limited.
@@
The penultimate row of the table shows the average RMSE obtained on all the data sets (purely indicative), confirming the interest of the proposed method for three of the five regressors. This average should be treated with caution. As each dataset is a problem of unique difficulty, the error scale differs between each dataset. So, if the new RMSE mean in test is higher, this does not necessarily mean that the regressors perform worse on average. For this reason, in the next section we present another view of the results for further analysis.
$\bullet$ Critical diagram - We present in Figure 5 critical diagrams comparing the results obtained by each regressor through the Nemenyi post-hoc test, performed after a Friedman signed rank test of RMSE values, taking into account all the data sets.
@@
From the graph on the left of Figure 5 it can be seen that the XGBoost regressor performs best, achieving the highest mean RMSE rank, closely followed by Random Forest (RF). These two regressors outperform the other three regressors by a wide margin, with linear regression being the worst performing model. It is closely followed by the naive Bayes regressor and the regression tree.The figure 5 in the middle compares the performance of all the regressors, with and without the proposed method: the first point we can observe is that, statistically, all the regressors with $X^{\prime}$ have a better average rank than their counterparts with only $X$. This is an encouraging result: no regressor has been negatively affected by the proposed method, since all regressors that have used this method are better ranked than their basic versions. Again, there are different groups, and the reader can compare the change in the ranking of different regressors by comparing the figure on the left with the one in the middle. Finally, the graph on the right of figure 5 shows that: when the proposed method is used, the regressors are no longer differentiable according to the Nemenyi test (only one group compared to the left figure), even though they have different mean ranks.